The diffrence in prices when a comodity is sold at a profit of 4% and the profit of 6% is Rs 3 find the selling price of of the comodity in the two cases?

Grade:8

1 Answers

Arun
6 years ago
Dear student
 
Let Cost = x
Then
At 6 % profit = x + 6x/100 = 106 x/100
At 4% profit = x + 4x/100 = 104 x/100
Now
(106x -104 x)/100  = 3
2x = 300
x = 150
Hence selling price at 6 % profit = 106 * 150/100 = 159 Rs
Selling price at 4 % profit = 104 * 150/100 = 156 Rs.
 
Regards
Arun (askIITians forum expert)